OpenStudy (anonymous):

PLEASE PLEASE PLEASE HELP!!!!! Which of the following describes ta cause of the flowering of literature and art by woman at the Heian (Kyoto) court around the year 1000 CE? -Woman educated in Japanese language, not chinese -Increasing presence of highly cultured courtesans -Men's concentration on military pursuits -Greater political influence of upper-class woman

OpenStudy (anonymous):

*A

OpenStudy (charlotte123):

Parmesan

OpenStudy (charlotte123):

Remember me, Parmesan Remember me

OpenStudy (charlotte123):

<3 -Char2014

OpenStudy (the_fizicx99):

She ate you at one point in her life

OpenStudy (anonymous):

parmesan, i can help if you help me

OpenStudy (anonymous):

i'm in need of a certain service

OpenStudy (anonymous):

do you comply?

OpenStudy (anonymous):

@eviI I will help you in any way I can. @charlotte123 im sorry but I dont know you.

OpenStudy (the_fizicx99):

So Char. ^ Idk why you didn't ask in your main ~

OpenStudy (charlotte123):

I messed up ;-;

OpenStudy (the_fizicx99):

Idiot, smh

OpenStudy (charlotte123):

Maybe It's because Parmesan Isn't me?

OpenStudy (tester97):

She knows you from austasia land @Parmesan

OpenStudy (anonymous):

thank you, the correct answer is A.

OpenStudy (charlotte123):

LOLOLOLOLOLOLOLOLOLOLO

OpenStudy (the_fizicx99):

No, it's C ~

OpenStudy (tester97):

No you idiots its D

OpenStudy (anonymous):

who do you trust? the purple haired boy, the weird not boy not girl person the pink haired clown, or me?

OpenStudy (charlotte123):

Idekrn

OpenStudy (anonymous):

Idk. .-.

OpenStudy (tester97):

Trust me. Natsu is always right

OpenStudy (anonymous):

Please be honest if you do or dont know the answer. .-.

OpenStudy (anonymous):

I do, it is the first choice.

OpenStudy (anonymous):

You can read this if you want proof.. http://en.wikipedia.org/wiki/Heian_period

OpenStudy (anonymous):

Alright well thank you! What do you need help with?

OpenStudy (tester97):

Dont trust @evil thats the wrong answer

OpenStudy (tester97):

Its D. D for wingspan

OpenStudy (charlotte123):

I'm dying Alley <3

OpenStudy (the_fizicx99):

No, it's A ~ dat retricecan't lie

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

Yw?

OpenStudy (tester97):

wOw i dont get a thank you? hella r00d

OpenStudy (anonymous):

parmeseana, did i not help :(

OpenStudy (charlotte123):

LOLOLOLOLOL

OpenStudy (the_fizicx99):

No, go away ^

OpenStudy (charlotte123):

no

OpenStudy (the_fizicx99):

Wrong arrow ;-;

OpenStudy (anonymous):

You did I said thank you to you too up there ^

OpenStudy (charlotte123):

tsk tsk tsk

OpenStudy (anonymous):

Parmeseana, i need help with solving this math problem: \[\large \int\limits_{0}^{\frac{ \pi }{ 2 }} \frac{ 1 }{ (1+x^2)(1+\tan~x) } dx = \frac{ \sqrt{\ln~\alpha} }{ 2 \ln^2~\beta}\] \[\text{Calculate} ~~\alpha~~~~ \text{and}~~~~~ \beta.\]

OpenStudy (the_fizicx99):

The answer is B, I need your help now

OpenStudy (anonymous):

that's alpha, and beta ( greek symbols).

OpenStudy (the_fizicx99):

;_; I noticed >.>

OpenStudy (anonymous):

@evil I have never seen anything like that before but I will try.